OpenStudy (cenaida):

Solve for x: (2x - 4)2 = 1 A. 2, 3 B. 3/2, 5/2 C. 3/2, 3/5 D. 3/2, -5/2

OpenStudy (anonymous):

first you disturb the 2

OpenStudy (anonymous):

or is it squared?? @cenaida

OpenStudy (cenaida):

its squared

OpenStudy (solomonzelman):

actually \[| 2x-4 | = 1\]

OpenStudy (anonymous):

i would do (2x-4)(2x-4)=1 so you have 2 answers for x

OpenStudy (anonymous):

so 4x^2-8x-8x+16

OpenStudy (solomonzelman):

\[| 2x-4 | = 1\]\[ 2x-4 = -1~~~~~~~~~and~~~~~~~~ 2x-4 = 1\]\[2x=3~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~2x=5\]\[x=3/2~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~x=5/2\]

OpenStudy (solomonzelman):

it's B

OpenStudy (solomonzelman):

Where did I get \[ | 2x-4 |=1\]LOOK. \[(2x-4)^2=1\]\[(2x-4)^2=1^2\]square both sides, the square is +- or an absolute value.

OpenStudy (solomonzelman):

I am sorry, but i suck at explaining, even teachers barely get my work....

OpenStudy (cenaida):

ok so its positive thank you

OpenStudy (solomonzelman):

Yes, it's positive. x= 3/2 and 5/2 See how I did it?